# Motivic Steenrod operations at the characteristic via infinite ramification

### v1

@@ 2.3. Motivic Cartan formula
(10) ×$\times\colon\mathcal{A}\otimes_{H\mathbb{F}^{\mathbb{A}^{1}}_{\ell}}\mathcal{A}\xrightarrow{\sim}\underline{\mathrm{maps}}(H\mathbb{F}^{\mathbb{A}^{1}}_{\ell}\otimes H\mathbb{F}^{\mathbb{A}^{1}}_{\ell},H\mathbb{F}^{\mathbb{A}^{1}}_{\ell})$
@@
In fact, more is true. Let $\Psi$ be the map (2) from the introduction, classifying admissible monomials in the dual $-\ell$ motivic Steenrod algebra
@@
Let $k$ be a field of characteristic $p\not=\ell$. Then, the following homotopies exist between maps to $\underline{\mathrm{maps}}(H\mathbb{F}^{\mathbb{A}^{1}}_{\ell}\otimes H\mathbb{F}^{\mathbb{A}^{1}}_{\ell},H\mathbb{F}^{\mathbb{A}^{1}}_{\ell})\in\mathrm{SH}_{k}$:
@@
    where $\tau$ is the non-zero element of $H_{\mathbb{A}^{1}}^{0}(k;\mathbb{F}_{2}(1))=\mu_{2}(k)$, and $\rho$ is the image of $-1$ in $H_{\mathbb{A}^{1}}^{1}(k;\mathbb{F}_{2}(1))=k^{\times}/(k^{\times})^{2}$.
@@ 3.1. A motivic Nizioł theorem
    $A$ is a discrete valuation ring of mixed characteristic $(0,p)$ with uniformizer $\pi$;
@@
(13) ${E(X)(n)}$${E(U,n)}$${E(X^{\prime})(n)}$${E(U^{\prime})(n)}$${E(X)(n)}$${E(U)(n).}$$\scriptstyle{p^{*}}$$\scriptstyle{j^{*}}$$\scriptstyle{r^{*}}$$\scriptstyle{p_{!}}$$\scriptstyle{j^{\prime*}}$$\scriptstyle{r_{!}}$$\scriptstyle{j^{*}}$
@@
[…]
@@
[…]
@@
[…]
@@
Variant 3.4 and Corollary 3.2 are motivic refinements of [AMM22, Corollary 3.3] and our methods are inspired by theirs. Unfortunately since absolute purity is not known more generally for motivic cohomology, we are unable to use the same method to prove a more general statement. Nevertheless, upcoming work of Bouis–Kundu [BK], however, has established an isomorphism:
@@ 3.2. Construction of power operations
    if $y\in H^{j}_{\mathbb{A}^{1}}(\mathfrak{X};\mathbb{F}_{p}(k))$ is such that $j-k<i$ and $k\leq i$, then $\mathrm{P}^{i}(y)=0$;
@@ 4.1. Non smoothable cycles at the characteristic
In this section, we give examples of algebraic cycles that cannot be smoothed modulo $p$, where $p$ is the characteristic of the base field $k$. The case of mod-$l$ coefficients was treated in [HRT74] using étale/singular cohomology. Let $R$ be ring of coefficients. The problem addressed in op. cit. was posed by Borel–Haefliger (for singular cohomology) [BH61]: when can write one cycle $\alpha\in R\otimes\mathrm{CH}^{n}(X)$ as a $R$-linear combination of nonsingular subvarieties of $X$? If $R=\mathbb{Z}$ then [HRT74, Theorem 1] gives a counterexample.
Our goal in this section is to give a counterexample to an even weaker version of the question in characteristic $p$ (so it has a better chance of being true): instead of asking that the subvariety is nonsingular we can ask that it is regularly immersed in $X$ [^8] Note that any smooth subvariety is regularly embedded.. In fact, the language of derived algebraic geometry [^9] We use very little of this theory; the formalism used in [KR18] more than suffices. provides a reasonably large class of immersions which we could hope to generate the group $R\otimes\mathrm{CH}^{n}(X)$.
@@
From the rest of this subsection, we fix $X$ to be a smooth $k$-scheme. If $Z\subset X$ is a closed subvariety such that $Z$ is smooth then we can associate to $Z$, its fundamental class $[Z]\in\mathrm{CH}^{\mathrm{codim}(Z)}(X)$. The work of Khan extends this to quasi-smooth closed immersions. In what follows, we use Khan’s extension of $\mathrm{SH}$ to derived schemes [Kha19a, Kha16]; note that by [Kha19a, Corollary 3.2.7] we have that $\mathrm{SH}_{X}\simeq\mathrm{SH}_{X^{\mathrm{cl}}}$ where $X^{\mathrm{cl}}\hookrightarrow X$ is the classical reduction of $X$. On $\mathrm{SH}_{X}$, the motivic spectrum $H\mathbb{Z}^{\mathbb{A}^{1}}$ is defined by pullback from $\mathrm{Spec}(\mathbb{Z})$.
@@
$[Z]:=\mathrm{tr}_{i*}(t(\mathcal{N}_{i}))\in\mathrm{CH}^{c}(X).$
@@
By [Kha19b, §3.3], the classes $[Z]\in\mathbb{Q}\otimes\mathrm{CH}^{\star}(X)$ coincide with the pushforwards of the virtual fundamental classes defined by Behrend–Fantechi [BF97]. Therefore, one can think of the construction in the above paragraph as an integral refinement of virtual fundamental classes. Moreover, $[Z]$ coincides with the usual fundamental class of $Z$ whenever $Z$ is regularly immersed
@@
[…]
@@
    for some cohomology class $c$. Thus, if $[Z]=\mathrm{tr}_{i*}(t)$, then $\mathrm{P}^{1}([Z])^{2}$ is divisible by $[Z]$. We show that this is impossible unless $\beta=0$ under certain assumptions on $m$ and $n$ that we are going to explain next.
    Let $m\geq 4$ be an even integer such that $-(m-2)/2$ is a non-square unit in $\mathbb{F}_{p}$, and let $n\geq m+2p+2$. Under the assumption on $n$, the cohomology ring of $X$ coincides with the polynomial ring generated by $c_{1},\dots,c_{m}$ in degrees less than equal to $22p+$ 2 [HRT74, §2]. Thus, it suffices to show that $\alpha c_{1}^{2}+\beta c_{2}$ does not divide $P^{1}(\alpha c_{1}^{2}+\beta c_{2})^{2}$ in $\mathbb{F}_{p}[c_{1},\dots,c_{m}]$. Furthermore, we may use the splitting principle to check this indivisibility in the ring $\mathbb{F}_{p}[a_{1},\dots,a_{m}]$, where $a_{i}$ are the Chern roots of $\mathcal{E}$. Lastly, we may assume that $\beta=1$.
@@
    By the assumption that $-(m-2)/2$ is not a square in $\mathbb{F}_{p}$, we have that $\xi_{\pm}^{p-1}=-1$, and therefore $\xi^{p+1}_{\pm}=\tfrac{m-2}{2}$. In particular $\xi_{\pm}$ are not roots of $g_{3}$, and thus $f_{3}$ does not divide the square of $g_{3}$, finishing the proof that a cohomology class of form $\alpha c_{1}+c_{2}$ is not quasi-smoothable. ∎
@@ 4.2. Obstructions against lifting to $\mathrm{MGL}$ and motivic Steenrod problem at the characteristic
[…]
@@ 4.3. The integral Tate conjecture at the characteristic
In this section, we formulate an optimist’s integral version of the crystalline Tate conjecture. Actually, we will immediately supply a counterexample to this conjecture so we should actually call it the integral Tate condition: the veracity of the statement is then an interesting, geometric condition on the variety in question. This version of the Tate conjecture does not seem to be as popular as its $\ell$-adic counterpart. To the authors’ knowledge, the only paper that discusses the rational version extensively is Morrow’s variational take on it [Mor19]. To formulate this conjecture, we recast the cycle class map via étale realization in the environment of non-$\mathbb{A}^{1}$-invariant motivic spectra $\mathrm{MS}_{S}$.
For the rest of the paper, $(\mathrm{Sm}^{?}_{X},\mathrm{MS}^{?}_{X})$ denotes either
@@
Let $S$ be a spectrum of a Dedekind domain or a field. Then
$(L_{\mathrm{et}}H\mathbb{Z}^{\mathbb{A}^{1}})^{{\kern-0.5pt}\wedge}_{p}=H\mathbb{Z}^{\mathrm{syn}}_{p}\in\mathrm{MS}_{S}.$
@@
We now observe that this map is 1) is an equivalence (essentially [Gei04, Theorem 1.2(2)]), 2) the first chern class (induced from the one coming from $H\mathbb{Z}^{\mathbb{A}^{1}}$) induces a $\mathbb{P}^{1}$-bundle formula on mod-$p^{r}$ syntomic cohomology and 3) mod-$p^{r}$ syntomic cohomology enjoys elementary blowup excision as verified in, say, [BEM, Section 4]. A combination of these three facts proves that we have an equivalence in $\mathrm{MS}_{S}$:
@@
For an equicharacteristic scheme $X$, the motivic cohomology of [EM23] defines a motivic spectrum $H\mathbb{Z}\in\mathrm{MS}_{X}$ since the multiplicative graded collection of presheaves $\{\mathbb{Z}(\star)^{\mathrm{mot}}\}$ has Nisnevich descent, has the $\mathbb{P}^{1}$-bundle formula (after a construction of a $c_{1}$) and enjoys elementary blowup excision [EM23, Theorem 1.1]. If $X$ is a field or, more generally, a regular noetherian equicharacteristic scheme then $H\mathbb{Z}\simeq H\mathbb{Z}^{\mathbb{A}^{1}}$ by [EM23, Theorem 6.1]. In this case, Proposition 4.13 reads: $(L_{\mathrm{et}}H\mathbb{Z})^{{\kern-0.5pt}\wedge}_{p}=H\mathbb{Z}^{\mathrm{syn}}_{p}$. For a more general scheme $S$, Bouis [Bou24] has constructed $H\mathbb{Z}\in\mathrm{MS}_{X}$ which extends the construction of [EM23]. In upcoming work of Bouis and Kundu, they have also established $H\mathbb{Z}\simeq H\mathbb{Z}^{\mathbb{A}^{1}}\in\mathrm{MS}_{X}$ for any Dedekind domain $S$. We expect (but cannot yet prove) that $H\mathbb{Z}\simeq H\mathbb{Z}^{\mathbb{A}^{1}}$ for a regular noetherian scheme $S$.
@@
We will be mostly interested in the case of $S=\mathrm{Spec}(k)$, a field of characteristic $p>0$.
@@
$[\bigoplus_{x\in X^{(n-1)}}W_{r}\Omega^{1}_{\log,\kappa(x)}\rightarrow\bigoplus_{x\in X^{(n)}}W_{r}\Omega^{0}_{\log,\kappa(x)}\rightarrow 0]$
@@
Remark 4.21 (Tate cycles).
Let $k$ be a perfect field of characteristic $p>0$ and $X$ a smooth, projective, geometrically connected $k$-variety. We define the integral crystalline Tate cycles to be the subgroup
$\mathrm{CrysTate}_{n}(X)=:\mathrm{Im}(H^{2n}_{\mathrm{syn}}(X;\mathbb{Z}_{p}(n))\rightarrow H^{2n}_{\mathrm{crys}}(X/W(k)))\subset H^{2n}_{\mathrm{crys}}(X/W(k))$
We remark that if $k$ is either a finite or algebraically closed field, then we have isomorphisms
$H^{2n}_{\mathrm{syn}}(X;\mathbb{Z}_{p}(n))_{\mathbb{Q}}\xrightarrow{\cong}\mathrm{CrysTate}_{n}(X)_{\mathbb{Q}}\cong(H^{2n}_{\mathrm{crys}}(X/W(k))[\tfrac{1}{p}])^{\phi=p^{n}};$
by [Mor19, Proposition 3.3(ii)]. This is the sense in which we are justified in calling these cycles to be an integral version of Tate cycles since the group $(H^{2n}_{\mathrm{crys}}(X/W(k))[\tfrac{1}{p}])^{\phi=p^{n}}$ is what is usually called the crystalline Tate cycles in the literature.
The crystalline cycle class map allows us to formulate the integral crystalline Tate conjecture (more rightly called, condition). Just as discussed by Schoen in [Sch98] there are a couple of different variants of this statement that one must be somewhat careful about; we stick to formulating these statements over finite fields in this paper. The integral crystalline Tate conjecture should be thought of as the $p$-adic analog of the integral $\ell$-adic Tate conjecture. The first systematic study of this conjecture and its counterexamples was first produced carefully in [CTS10]. Other counterexamples to the $\ell$-adic integral Tate conjecture have been studied by various authors including (but not limited to): [PY15, Ant16, Qui11, Ben25].
@@
Let $k$ be a finite field and $\overline{k}$ be an algebraic closure. Let $X$ be a smooth, projective, geometrically connected $k$-variety and let $n\geq 0$ be fixed. Consider the maps
(22) $\mathrm{cyc}^{n}_{\mathrm{crys}}:\mathrm{CH}^{n}(X)\otimes\mathbb{Z}_{p}\rightarrow\mathrm{CrysTate}_{n}(X).$
(23) $\mathrm{cyc}^{n}_{\mathrm{crys}}:\mathrm{CH}^{n}(X_{\overline{k}})\otimes\mathbb{Z}_{p}\cong\mathrm{colim}_{k\subset k^{\prime}\subset\overline{k}}\mathrm{CH}^{n}(X_{k^{\prime}})\otimes\mathbb{Z}_{p}\rightarrow\mathrm{colim}_{k\subset k^{\prime}\subset\overline{k}}\mathrm{CrysTate}_{n}(X_{k^{\prime}}).$
where the colimit is taken along all finite extensions of $k$ inside $\overline{k}$. We consider the following statements:
  1. $(T^{n}(X;\mathbb{Z}_{p}))$
    The map (22) is surjective.
  2. $(T_{\mathrm{cts}}^{n}(X_{\overline{k}};\mathbb{Z}_{p}))$
    The map (23) is surjective.

### v2

@@ 2.3. Motivic Cartan formula
(10) $\times\colon\mathcal{A}\otimes_{H\mathbb{F}^{\mathbb{A}^{1}}_{\ell}}\mathcal{A}\xrightarrow{\sim}\underline{\mathrm{maps}}(H\mathbb{F}^{\mathbb{A}^{1}}_{\ell}\otimes H\mathbb{F}^{\mathbb{A}^{1}}_{\ell},H\mathbb{F}^{\mathbb{A}^{1}}_{\ell})$
@@
In fact, more is true. Let $\Psi$ be the map (2) from the introduction, classifying admissible monomials in the dual mod-$\ell$ motivic Steenrod algebra
@@
Let $k$ be a field of characteristic $p\not$$\ell$. Then, the following homotopies exist between maps to $\underline{\mathrm{maps}}(H\mathbb{F}^{\mathbb{A}^{1}}_{\ell}\otimes H\mathbb{F}^{\mathbb{A}^{1}}_{\ell},H\mathbb{F}^{\mathbb{A}^{1}}_{\ell})\in\mathrm{SH}_{k}$:
@@
    where $\tau$ is the non-zero element of $H_{\mathbb{A}^{1}}^{0}(k;\mathbb{F}_{2}(1))=\mu_{2}(k)$, and $\rho$ is the image of 1-1 in $H_{\mathbb{A}^{1}}^{1}(k;\mathbb{F}_{2}(1))=k^{\times}/(k^{\times})^{2}$.
@@ 3.1. A motivic Nizioł theorem
    AA is a discrete valuation ring of mixed characteristic $(0,p)$ with uniformizer $\pi$;
@@
[…]
@@
We now show that $p_{!}(1_{X^{\prime}})\simeq d$. By assumption, $p$ factors as $X^{\prime}\stackrel{{\scriptstyle\iota}}{{\hookrightarrow}}\mathbb{P}^{1}_{X}\xrightarrow{\pi}X$, where $\iota$ is a divisor for the line bundle $\mathcal{O}(d)$, and $\pi$ is the projection. As the second power of $c_{1}(\mathcal{O}(1))$ vanishes in $E$-cohomology, we can use the formal group law of $E$ to compute that $\iota_{!}(1_{X^{\prime}})=dc_{1}(\mathcal{O}(1))$. As $\pi_{!}(c_{1}(\mathcal{O}(1)))=1_{X}$, it follows that $p_{!}(1_{X^{\prime}})=d$ in $E$-cohomology, as desired.
@@
$p_{*}p^{*}$$p_{*}k_{*}k^{!}p^{*}$$\Sigma^{-2,-1}p_{*}k_{*}k^{*}p^{*}$$p_{*}p^{!}$$p_{*}k_{*}k^{!}p^{!}$$i_{*}q_{*}q^{!}i^{!}$$i_{*}q_{*}q^{*}i^{*}$$\mathrm{id}$$i_{*}i^{!}$$\scriptstyle{\mathfrak{p}_{p}}$$\scriptstyle{\eta}$$\scriptstyle{\mathfrak{p}_{p}}$$\scriptstyle{\mathfrak{p}_{k},\simeq}$$\scriptstyle{\mathfrak{p}_{p\circ k}}\scriptstyle{\simeq}$$\scriptstyle{\eta}$$\scriptstyle{\eta}\scriptstyle{\simeq}$$\scriptstyle{\eta,\simeq}$$\scriptstyle{\mathfrak{p}_{i\circ q}}\scriptstyle{\simeq}$$\scriptstyle{\eta}$
@@
$Y^{\prime}_{mm^{\prime}}$$X_{mm^{\prime}}$$Y_{m}$$X_{m}.$$\scriptstyle{q}$$\scriptstyle{k}$$\scriptstyle{p}$$\scriptstyle{i}$
@@
Variant 3.4 and Corollary 3.2 are motivic refinements of [AMM22, Corollary 3.3] and our methods are inspired by theirs. Unfortunately since absolute purity is not known more generally for motivic cohomology, we are unable to use the same method to prove a more general statement. Nevertheless, the work of Bouis–Kundu [BK25], however, has established an isomorphism:
@@ 3.2. Construction of power operations
    if $y\in H^{j}_{\mathbb{A}^{1}}(\mathfrak{X};\mathbb{F}_{p}(k))$ is such that jk<ij-k<i and $k\leq i$, then $\mathrm{P}^{i}(y)=0$;
@@ 4.1. Non smoothable cycles at the characteristic
In this section, we give examples of algebraic cycles that cannot be smoothed modulo $p$, where $p$ is the characteristic of the base field $k$. The case of mod-$l$ coefficients was treated in [HRT74] using étale/singular cohomology.
Let $R$ be ring of coefficients. The problem addressed in op. cit. was posed by Borel–Haefliger (for singular cohomology) [BH61]: when can write one cycle $\alpha\in R\otimes\mathrm{CH}^{n}(X)$ as a $R$-linear combination of nonsingular subvarieties of $X$? If $R=\mathbb{Z}$ then [HRT74, Theorem 1] gives a counterexample.
Our goal in this section is to give a counterexample to an even weaker version of the question in characteristic $p$ (so it has a better chance of being true): instead of asking that the subvariety is nonsingular we can ask that it is regularly immersed in XX [^9] Note that any smooth subvariety is regularly embedded.. In fact, the language of derived algebraic geometry [^10] We use very little of this theory; the formalism used in [KR18] more than suffices. provides a reasonably large class of immersions which we could hope to generate the group $R\otimes\mathrm{CH}^{n}(X)$.
@@
From the rest of this subsection, we fix $X$ to be a smooth $k$-scheme. If $Z\subset X$ is a closed subvariety such that $Z$ is smooth then we can associate to $Z$, its fundamental class $Z]\in\mathrm{CH}^{\mathrm{codim}(Z)}(X)$. The work of Khan extends this to quasi-smooth closed immersions. In what follows, we use Khan’s extension of $\mathrm{SH}$ to derived schemes [Kha19a, Kha16]; note that by [Kha19a, Corollary 3.2.7] we have that $\mathrm{SH}_{X}\simeq\mathrm{SH}_{X^{\mathrm{cl}}}$ where $X^{\mathrm{cl}}\hookrightarrow X$ is the classical reduction of $X$. On $\mathrm{SH}_{X}$, the motivic spectrum $H\mathbb{Z}^{\mathbb{A}^{1}}$ is defined by pullback from $\mathrm{Spec}(\mathbb{Z})$.
@@
$Z]:=\mathrm{tr}_{i*}(t(\mathcal{N}_{i}))\in\mathrm{CH}^{c}(X).$
@@
By [Kha19b, §3.3], the classes $Z]\in\mathbb{Q}\otimes\mathrm{CH}^{\star}(X)$ coincide with the pushforwards of the virtual fundamental classes defined by Behrend–Fantechi [BF97]. Therefore, one can think of the construction in the above paragraph as an integral refinement of virtual fundamental classes. Moreover, $[Z]$ coincides with the usual fundamental class of $Z$ whenever $Z$ is regularly immersed
@@
We remark on the compatibility of $\mathrm{tr}_{i*}$ and power operations. A motivic cohomology operation on $H\mathbb{F}_{p}^{\mathbb{A}^{1}}$ is a morphism in $\mathrm{SH}_{X}$ given by $H\mathbb{F}_{p}^{\mathbb{A}^{1}}\xrightarrow{o}\Sigma^{n}_{\mathbb{T}}H\mathbb{F}_{p}^{\mathbb{A}^{1}}[m]$. Since the formation of $\mathrm{tr}_{i}$ upgrades to a transformation
@@
    for some cohomology class $c$. Thus, if $Z]=\mathrm{tr}_{i*}(t)$, then $\mathrm{P}^{1}([Z])^{2}$ is divisible by $[Z]$. We show that this is impossible unless $\beta=0$ under certain assumptions on $m$ and $n$ that we are going to explain next.
    Let $m\geq 4$ be an even integer such that (m2)/2-(m-2)/2 is a non-square unit in $\mathbb{F}_{p}$, and let $n\geq m+2p+2$. Under the assumption on $n$, the cohomology ring of $X$ coincides with the polynomial ring generated by $c_{1},\dots,c_{m}$ in degrees less than equal to $2p+2$ [HRT74, §2]. Thus, it suffices to show that $\alpha c_{1}^{2}+\beta c_{2}$ does not divide $P^{1}(\alpha c_{1}^{2}+\beta c_{2})^{2}$ in $\mathbb{F}_{p}[c_{1},\dots,c_{m}]$. Furthermore, we may use the splitting principle to check this indivisibility in the ring $\mathbb{F}_{p}[a_{1},\dots,a_{m}]$, where $a_{i}$ are the Chern roots of $\mathcal{E}$. Lastly, we may assume that $\beta=1$.
@@
    By the assumption that (m2)/2-(m-2)/2 is not a square in $\mathbb{F}_{p}$, we have that $\xi_{\pm}^{p-1}=-1$, and therefore $\xi^{p+1}_{\pm}=\tfrac{m-2}{2}$. In particular $\xi_{\pm}$ are not roots of $g_{3}$, and thus $f_{3}$ does not divide the square of $g_{3}$, finishing the proof that a cohomology class of form $\alpha c_{1}+c_{2}$ is not quasi-smoothable. ∎
@@ 4.2. Obstructions against lifting to $\mathrm{MGL}$ and motivic Steenrod problem at the characteristic
    $\underline{\Omega}^{n}(X)_{(p)}$$\pi_{-2n}\mathrm{MGL}(n)(X)_{(p)}$$H^{2n}_{\mathrm{mot}}(X;\mathbb{Z}_{(p)}(n))$$H^{2n}_{\mathbb{A}^{1}}(X;\mathbb{Z}_{(p)}(n)),$}
@@ 4.3. The integral Tate conjecture at the characteristic
In this section, we formulate optimistic versions of the crystalline Tate conjecture. Actually, we will immediately supply a counterexample to this conjecture so we should actually call it the integral Tate condition: the veracity of the statement is then an interesting, geometric condition on the variety in question. This version of the Tate conjecture does not seem to be as popular as its $\ell$-adic counterpart.
Remark 4.13.
To the authors’ knowledge, the first paper that discusses the rational version extensively is Morrow’s variational take on it [Mor19]. An integral variant for divisors, using flat cohomology for the target, was first studied by D’Addezio [D’A24] who proved it for abelian varieties over finite fields.
[…]
@@
Let $S$ be a spectrum of a Dedekind domain or a field. Then there is an equivalence of $\mathbb{E}_{\infty}$-algebras in $\mathrm{MS}_{S}$:
$(L_{\mathrm{et}}H\mathbb{Z}^{\mathbb{A}^{1}})^{{\kern-0.5pt}\wedge}_{p}\xrightarrow{\simeq}H\mathbb{Z}^{\mathrm{syn}}_{p}\in\mathrm{MS}_{S}.$
@@
We now observe that this map is 1) is an equivalence (essentially [Gei04, Theorem 1.2(2)]), 2) the first chern class (induced from the one coming from $H\mathbb{Z}^{\mathbb{A}^{1}}$) induces a $\mathbb{P}^{1}$-bundle formula on mod-$p^{r}$ syntomic cohomology and 3) mod-$p^{r}$ syntomic cohomology enjoys elementary blowup excision as verified in, say, [BEM, Section 4]. A combination of these three facts proves that we have a multiplicative equivalence in $\mathrm{MS}_{S}$:
@@
For an equicharacteristic scheme $X$, the motivic cohomology of [EM23] defines a motivic spectrum $H\mathbb{Z}\in\mathrm{MS}_{X}$ since the multiplicative graded collection of presheaves $\{\mathbb{Z}(\star)^{\mathrm{mot}}\}$ has Nisnevich descent, has the $\mathbb{P}^{1}$-bundle formula (after a construction of a $c_{1}$) and enjoys elementary blowup excision [EM23, Theorem 1.1]. If $X$ is a field or, more generally, a regular noetherian equicharacteristic scheme then $H\mathbb{Z}\simeq H\mathbb{Z}^{\mathbb{A}^{1}}$ by [EM23, Theorem 6.1]. In this case, Proposition 4.14 reads:
(20) $(L_{\mathrm{et}}H\mathbb{Z})^{{\kern-0.5pt}\wedge}_{p}=H\mathbb{Z}^{\mathrm{syn}}_{p}.$
For a more general scheme $S$, Bouis [Bou24] has constructed $H\mathbb{Z}\in\mathrm{MS}_{X}$ which extends the construction of [EM23]. In upcoming work of Bouis and Kundu, they have also established $H\mathbb{Z}\simeq H\mathbb{Z}^{\mathbb{A}^{1}}\in\mathrm{MS}_{X}$ for any Dedekind domain $S$. We expect (but cannot yet prove) that $H\mathbb{Z}\simeq H\mathbb{Z}^{\mathbb{A}^{1}}$ for a regular noetherian scheme $S$.
@@
where the equivalence is (20).
We will be mostly interested in the case of $S=\mathrm{Spec}(k)$, a perfect field of characteristic $p>0$.
@@
$\bigoplus_{x\in X^{(n-1)}}W_{r}\Omega^{1}_{\log,\kappa(x)}\rightarrow\bigoplus_{x\in X^{(n)}}W_{r}\Omega^{0}_{\log,\kappa(x)}\rightarrow 0]$
@@
Definition 4.22 (Tate cycles).
Let $k$ be a perfect field of characteristic $p>0$ and $X$ a smooth, projective $k$-variety. For an extension $k^{\prime}/k$ we write $G_{k^{\prime}/k}:=\mathrm{Gal}(\overline{k}/k)$; if $k^{\prime}$ is an algebraic closure of $k$ then write $G_{k}:=G_{k^{\prime}/k}$
  1. (1)
    The abelian group of integral $p$-adic Tate cycles of codimension $n$ on $X$ is given by
    $\mathrm{Tate}^{n}_{p}(X):=\mathrm{Im}(H^{2n}_{\mathrm{syn}}(X,\mathbb{Z}_{p}(n))\rightarrow H^{2n}_{\mathrm{crys}}(X/W(k))))$
  2. (2)
    Let $k^{\prime}/k$ be an extension of $k$ (possibly infinite, like an algebraic closure). We define:
    $\mathrm{Tate}^{n}_{p}(X_{k^{\prime}})^{G_{k^{\prime}/k}}=:\mathrm{Im}(H^{2n}_{\mathrm{syn}}(X_{k^{\prime}},\mathbb{Z}_{p}(n))^{G_{k^{\prime}/k}}\rightarrow H^{2n}_{\mathrm{crys}}(X_{k^{\prime}}/W(k))^{G_{k^{\prime}/k}})$
Remark 4.23 (The case of finite fields).
Let $k$ be a finite field and $\overline{k}$ be an algebraic closure. Then the Hochschild–Serre spectral sequence for $\widehat{\mathbb{Z}}\xrightarrow{\cong}G_{k}$:
$H^{r}(\widehat{\mathbb{Z}},H_{\mathrm{syn}}^{s}(X_{\overline{k}},\mathbb{Z}_{p}(j)))\Rightarrow H^{r+s}_{\mathrm{syn}}(X,\mathbb{Z}_{p}(j))\qquad\forall j\geq 0,$
degenerates (since $\widehat{\mathbb{Z}}$ has $p$-cohomological dimension one) to produce a short exact sequence
$0\rightarrow H^{1}(\widehat{\mathbb{Z}},H^{i-1}_{\mathrm{syn}}(X_{\overline{k}},\mathbb{Z}_{p}(j)))\rightarrow H_{\mathrm{syn}}^{i}(X,\mathbb{Z}_{p}(j))\rightarrow H_{\mathrm{syn}}^{i}(X_{\overline{k}},\mathbb{Z}_{p}(j))^{\widehat{\mathbb{Z}}}\rightarrow 0$
In particular, unwinding Definition 4.22, we get a surjection
(23) $\mathrm{Tate}^{j}_{p}(X)\twoheadrightarrow\mathrm{Tate}^{j}_{p}(X_{\overline{k}})^{G_{k}}$
Remark 4.24 (Relationship with rationalized Tate cycles).
$k$ is either a finite or algebraically closed field, then we have isomorphisms
(24) $H^{2n}_{\mathrm{syn}}(X;\mathbb{Z}_{p}(n))\otimes{\mathbb{Q}}\xrightarrow{\cong}\mathrm{Tate}^{n}_{p}(X)\otimes{\mathbb{Q}}\cong(H^{2n}_{\mathrm{crys}}(X/W(k))[\tfrac{1}{p}])^{\phi=p^{n}};$
by [Mor19, Proposition 3.3(ii)]. This is the sense in which we are justified in calling these cycles to be an integral version of Tate cycles since the group $(H^{2n}_{\mathrm{crys}}(X/W(k))[\tfrac{1}{p}])^{\phi=p^{n}}$ is what is usually called the crystalline version of Tate cycles in the literature.
The crystalline cycle class map allows us to formulate the integral crystalline Tate conjecture (more rightly called, condition). Just as discussed by Schoen in [Sch98] there are a couple of different variants of this statement that one must be somewhat careful about; we stick to formulating these statements over finite fields in this paper. The integral crystalline Tate conjecture should be thought of as the $p$-adic analog of the integral $\ell$-adic Tate conjecture.
For the rest of the section, by the cycle class map we mean the composition of the first two maps in the sequence
(25) $\mathrm{CH}^{n}(X)\otimes\mathbb{Z}_{p}\to H^{2n}_{\mathrm{syn}}(X,\mathbb{Z}_{p}(n))\twoheadrightarrow\mathrm{Tate}^{n}_{p}(X)\subset H^{2n}_{\mathrm{crys}}(X/W(k)),$
and it is denoted by $\mathrm{cyc}^{n}_{X}$.
@@
Let $k$ be a finite field and let $X$ be a smooth, projective, geometrically connected $k$-variety. For any $j\geq 0$, we define the following conditions:
  1. $(A^{n}_{X_{k}})$
    The cycle class map
    $\mathrm{cyc}_{X}^{n}:\mathrm{CH}^{n}(X)\otimes\mathbb{Z}_{p}\rightarrow\mathrm{Tate}^{n}_{p}(X)$
    is surjective.
  2. $(B^{n}_{X_{k}})$
    The cycle class map composed with (23)
    $\overline{\mathrm{cyc}}_{X}^{n}:\mathrm{CH}^{n}(X)\otimes\mathbb{Z}_{p}\rightarrow\mathrm{Tate}_{p}^{n}(X_{\overline{k}})^{G_{k}}$
    is surjective.
  3. $(C^{n}_{X_{k}})$
    The colimit of the system of maps
    $\overline{\mathrm{cyc}}_{X_{k^{\prime}}}^{n}:\mathrm{CH}^{n}(X_{k^{\prime}})\otimes\mathbb{Z}_{p}\rightarrow\mathrm{Tate}^{n}_{p}(X_{\overline{k}})^{G_{k^{\prime}}}\qquad\text{$k^{\prime}/k$ finite extension},$
    which defines the continuous cycle class map
    $\mathrm{cyc}_{\mathrm{cts},X}^{n}:\mathrm{CH}^{n}(X_{\overline{k}})\otimes\mathbb{Z}_{p}\rightarrow\mathrm{Im}(\mathrm{colim}_{k\subset k^{\prime}\subset\overline{k}}\mathrm{Tate}_{p}^{n}(X)^{G_{k^{\prime}}}\rightarrow H^{2n}_{\mathrm{crys}}(X_{\overline{k}}/W(k)))$
    is surjective.